OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6

OpenStudy (anonymous):

looking for x?

OpenStudy (anonymous):

@Goldenmoon are you looking for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if u r then do the following: distribute: 3x+3=-2x+2+6 combine like terms: 5x=5 divide 5 from both sides: x=1

OpenStudy (anonymous):

for some reason my answer came out as -1 ;/

OpenStudy (anonymous):

no u did ur math wrong i think

OpenStudy (anonymous):

yeah , may you help with one more pretty please ?

OpenStudy (anonymous):

sure what is it?

OpenStudy (anonymous):

3 over 4 x + 5 over 8 = 4x

OpenStudy (anonymous):

so...